OpenStudy (anonymous):

A jar of dimes and quarters contains $15.25. There are 103 coins in all. How many quarters are in the jar?

OpenStudy (lncognlto):

Let the number of dimes = x and the number of quarters = y Then two equations that can be formed from this information are: x + y = 103 (the number of coins) and 10x + 25y = 1525 (the total value in cents) Now solving these equations, what do we get?

OpenStudy (anonymous):

i need the answer now or i fail highschool

OpenStudy (lncognlto):

>.> Well, first, what do you think the answer is?

OpenStudy (anonymous):

3.....

OpenStudy (anonymous):

I has no idea bruh .-.

OpenStudy (tester97):

how will you fail the entire year of high school just by missing one question? o.O

OpenStudy (anonymous):

im in credit recovery HELP!!!!!!!

OpenStudy (tester97):

AND on this site we dont give out answers we help guide the asker to the correct answer. So take some time to read the Coc here is the link: http://openstudy.com/code-of-conduct

OpenStudy (anonymous):

THATS A LOAD OF CRAP!!! CMON HELP ME

OpenStudy (tester97):

NO ITS NOT A LOAD OF CRAP it is the honest truth. Maybe you would know that if you read the Coc.

OpenStudy (anonymous):

I just reported you .-.

OpenStudy (anonymous):

I mean I reported retarded .-.

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

HUZZAH :P

OpenStudy (anonymous):

You keep cussing .-.

OpenStudy (anonymous):

f you mother licker ive got a mangina

OpenStudy (anonymous):

You dirty mouth :(

OpenStudy (anonymous):

*gets peaches*

OpenStudy (tester97):

*clicks report* now please follow the code of conduct it could really help you to be a better user on this site and make it more enjoyable for others. good day. ^.^

OpenStudy (anonymous):

I just reported twice .-.

OpenStudy (anonymous):

*high fives tester*

OpenStudy (anonymous):

@helpimretarded, YOU CUSS ONE MO TIME, AND YOU GETTIN MY PEACHES IN YO FACE :(

OpenStudy (anonymous):

i willl eat your peaches all up

OpenStudy (anonymous):

i willl eat your peaches all up

OpenStudy (anonymous):

That's the best you can do bruh?

OpenStudy (anonymous):

@tester97, get out the peaches .-.

OpenStudy (anonymous):

i have been known to do better

OpenStudy (anonymous):

Yeah, right.

OpenStudy (tester97):

ok ur wastin my time @helpimretarded so imma go i tried to b nice and everything so its time for me to go. good day!

OpenStudy (anonymous):

*leaves*